The atmospheric pressure at sea level is about 1 x 10^5 Pa. What is the force exerted on the top of a office desk, 152 cm long and 76 cm wide?
1
Expert's answer
2013-01-11T04:20:57-0500
We find area of the desk:
S = 152cm*76cm = 11552 cm^2 = 1.1552 m^2
Now we find force
F=p*S = 1 x 10^5 Pa * 1.1552 m^2 = 115 520 N